These are the news starting 01 Jan 2016
Click here for the news of the previous years.

31 Dec 16 Made overplot graph of Uccle and meteoLCD total ozone column measurements for 2016. See "DOBSON" page.
16 Dec 16 Recalibration of CO2 sensor. Zero was offset by 12 ppm. Multiple checks with zero and span. Seems ok now.
12 Dec 16 Recalibration and multiple checks on AC31m (NOx,NO).
10 Dec 16 1. Our very, very old and nearly defunct Xmas lighting has been replaced by a brand-new LED strip (50m wound around the handrail). Take a look if you are passing by at night-time!
2. The AC31M NOx sensor causes some serious headaches again:
2a. Controlling all connections, I saw that we registered as the NO2 signal what in reality was the NOx signal (NO2 = NOx - NO). Normally the NO2 levels should be lower than the NO. This mistake goes back for a very long time; I guess since the AC31M was installed. I will change the GNUPLOT graph to show the three NOx. NO and NO2 levels, and the channel 16 column of the logger data file will be re-labeled NOx. All previous yearly Excel files will be modified by re-labeling and adding a calculated NO2 column. The trends also must be changed. All this will take some time, so thank you for your patience
2b. I was unable to correctly recalibrate the AC31M to the level of our span gas. The shown levels could be too low. This b****y NOx sensor has so many dysfunctions that I really think on shutting down our NOx measurements.
08 Dec 16 1. Maintenance by T.S.E.
2: I changed the dessicant capsules of the UVB and UVA sensors (disregard spikes in data series).
02 Nov 16 There was a temporary problem with the half-hour transmission between the datalogger and its calling computer at 12:00 UTC. I am not sure why the communication failed, but manual intervention and a good reboot solved the problem.
01 Nov 16 Upload of oct16.dat into data archive.
30 Oct 16 The NO/NO2 sensor remained stuck in Zero mode since last maintenance from 21 October (my fault :-(( ).Correct operational mode restarted today at 07:50 UTC.
This long-time zero operation suggests about the same offsets to be subtracted from the readings as those given below: subtract 3 ug/m3 from the NO readings, and 10 ug/m3 from the NO2 readings.
21 Oct 16 Maintenance and calibration. The NO/NO2 measurements still have uncorrectable zero offfsets, which became sightly smaller as given in the 25 Sep news. Subtract 3 ug/m3 from the NO and 12 ug/m3 from the NO2 readings.
02 Oct 16 Upload of sep16.dat into data archive
25 Sep 16 1. There was a problem with lunar phases in the smartphone APP; this has been corrected by Marcel Kramer. Thx to Mr. Zimmer for detecting and reporting this problem.
2. Recalibration of the NO/NO2 sensor: there remains an uncorrectable problem with the zero offset. The NO readings (as displayed on the graphs and in the data) are about 4 ug/m3 too high, the NO2 readings 15 ug/m3 too high. Simply subtract 4 from the NO and 15 from the NO2 readings; the raw data and the plots will not be changed, and the plots will continue to represent the raw data.
3.Recalibration of the CO2 sensor.
4. I bought today on the flea-market hold in the premises of the LCD an UV-tester for checking banknotes. It comes handy as an UVA source for quick checks of our UVA Biometer and a small vintage J-221 photovoltaic UV intensity meter. Photos here, here (outside test) and here (very kludgy outside test, sensor protected from sunrays)
16 Sep 16 1. Maintenance work by T.S.E. The NOx sensor's main pump has been replaced, and it will run several days in test-mode to allow the molybden catalysator regain a normal state.
2. The dessicant capsules of the UVB and UVA sensors have been refreshed. The sensors were taken off-line for about 1.5 hours. Ignore the data from 16Sep16 08:30 to 09:30 UTC.
01 Sep 16 Upload of aug16.dat into data archive
15 Aug 16 1. As I presumed, the main pump was down, as well as the NOx sensor; the fault was a tripped fuse. I inspected the pump, which seems ok. Restart everything at 07:20 UTC. Will have to watch the future behavior of that pump, as it might be the culprit of the excess current drawn. Well, after half an hour, the sample indicator of the CO2 sensor gives a warning, so the pump was dismounted again and replaced by a new one (which by lucky insight we had among our spares!). Restart is at  08:20 UTC.
2. Recalibration & maintenance CO2 and O3. The pump of the NOx sensor is dead (yet another pump!), so no NO and NO2 until the next maintenance by T.S.E.
11 Aug 16 The black series continues..The ground ozone and CO2 levels are much too low, possibly due to a malfunction of the main pump. Repair is not possible before the 15th August.
added 16 Aug.: the external pump of the NO2 sensor tripped a fuse, which shut down the main pump. So all CO2 and O3 readings from 02Aug 15:30 to 15Aug should be ignored, as they are too low.
08 Aug 16 The NOx sensor is out of order since 02Aug2016 15:30. Please be patient, an eventual repair can not be made before 7 days.
01 Aug 16 Upload of jul16.dat into data archive. Missing Dobson measurements will follow asap.
13 Jul 16 Maintenance and calibration work.
12 Jul 16 A new calibration factor of 0.9422 is applied to the total ozone column Microtops #3012 readings (made by Mike Zimmer) starting 02Jun2016. A new intercomparison suggests this calibration factor which replaces the previous of 0.8822.
11 Jul 16 From 07:00 to 09:00: the temperature sensitive voltage amplifier of the main PT100 sensor (AIT, ch.2) has been replaced by an (expensive)  Phoenix Contact MCR model with a 4..20 mA output and zero and span adjustments potentiometers. Calibration and verifications have been made using a Bürstner PT100 simulator auctioned at eBay; this simulator is extremely handy if your amplifier is adjustable. The data of this morning may contain some glitches.
01 Jul 16 Uploaded June 2016 (jun16.dat) data into data archive.
21 Jun 16 Re-check calibration of AC31M and CO2 sensors. The AC31M (NO/NO2) still has an uncorrectable zero offset of about + 10 ug/m3 for NO2 and +5 ug/m3 for NO.
17 Jun
Maintenance by T.S.E.
11 Jun 16 The "Promotion 1968" (i.e. the students who made their final exam in 1968) met at the LCD, and it was a pleasure for me to give them a short exposé on climate change titled: "Changement climatique, pas de panique!" The Powerpoint presentation is here.
01 Jun 16 Uploaded May 2016 (may16.dat) data into data archive. Due to the change of logger  computer, there are missing data from 03 May 00:00 to 20:30 UTC.
24 May 16 RTL TV aired a short video on Al Gore's "An Inconvenient Truth" 10th anniversary. I make a very short comment at the end; the video showing some meteoLCD stuff and myself are from archiv material, as my comment was made through a phone line. See here (video courtesy RTL.LU). If your browser tells you that the file is corrupt and/or does not show any image, do not believe, but download the video and play it back with VLC or Quicktime.
18 May 16 1. Calibration and maintenance work. The zero bias of the NO and NO2 readings is confirmed.
2. Restarted the WBGT sensor (even if weather is far from heat-stressing).
09 May 16 1. The NO/NO2 sensors has a positive bias of NO: 1.9 ug/m3 and NO2: 10.6 ug/m3. It has been recalibrated at ~08:00, but remember that NO readings might be 2 ug/m3 too high, and NO2 readings 11 ug/m3 too high.

2. I had to make some work on the new HP Thinclients (main logger computer and Davis backup station computer) --> possible interruption of 1 hour.

3. Today was the day of the rare Mercury transit (planet Mercury passes in front of the solar disk). As clouds cleared up after 15:30 local time we (FM and CH) quickly set up an telescope and tried some pictures.

Some more: photo1, photo2, photo3 (I hope I got this right, and did not confuse Mercury and sunspot :-)).
See video and pictogram of path of the transit as seen from near-by Ettelbruck town. Our photo is among the collection shown at the website.

08 May 16 The NO/NO2 runs in zero mode for a couple of days, to get an idea of the zero offset and to reconfigure the molybdenum in the oven.
04 May 16 As predicted, there were some serious problems when switching to the new HP computer system and a new command-line batch processing; we started with a cryptic error message from the main command-line program. Finally the cause was found being a timing problem between the internal logger operation and the schedule for retrieving the stored data. Everything seems to be ok since this morning 05:00 UTC.
The PT100 sensor is now routed through an amplifier mounted close to the sensor on the terrace. The [0...10V] output (corresponding to [-50...100 °C] span will travel down to the main rack, and there goes through a voltage-divider as the logger has a maximum input of 2V. Reconfiguration of the sensor means that we had to stop and empty the logger. No data (at least not many) are lost....
03 May 16 Raoul and I will do some work on the PT100 sensor (we will add an amplifier) and on the logger computer (which will be replaced by an exlusively solid-state HP Thinclient running XPe). Be prepared for some interruption!
01 May 16 1. Uploaded April 2016 (apr16.dat) data into data archive.
2. NO/NO2 sensor again online (operated in forced zero mode for the last 48 hours after the maintenance by T.S.E. to stabilize the molybdenum in the oxidation chamber). Ignore the last 7 days values. NO2 readings seem to have a positive bias of about 11 ug/m3; use with care.
29 Apr 16 1. Repaired Davis Vantage Pro Plus solar radiation sensor. Humidity sensor is definitely dead, will not be replaced.
2. Maintenance by T.S.E. The NO/NO2 sensor will be go online probably next Tuesday.
18 Apr 16 1. Our lightning computer behaves erratically, and goes into non-responsive state after a couple of days. I replaced this morning the old faithful E600 FSC computer by an identical one, keeping the harddisk and all software from the unhealthy one.
2. I replaced the Weatherlink 5.1. software of the Davis backup station with version 5.6., which correctly identifies the Vantage Pro Plus model we are using (that is the original Vantage Pro with additional total solar and UVB sensors. The HP T5370 thin client seems to behave very well!
3. A group of  "Erasmus" students from various countries (Iceland, Italy...) visited meteoLCD. Prof. Geneviève Harles manages this group for a whole week; quite a job!
17 Apr 16 Some important changes are on the way. Since a couple of months the old faithful fanless SUMICOM computers which drive a lot of our sensors are starting to develop geriatric problems (they are about 8 years old, and have passive cooling but a traditional IDE hard-disk). I was looking for a completely passive (nothing rotating) and inexpensive replacement, and found a rather good solutions (in my opinion): the HP TT5370 thin client with embedded XP (XPe). These small computers have 1 GB Ram and 1 GB flash HD (the older DOC = disk on chip). About 400 MB are available for usage, after the system and a couple of programs are installed. That's not much, but enough for the single job to be done plus Automate scheduler plus TightVNC). What shines is that all software is still available at HP, and that cloning a machine on an USB stick is very easy (restoring is by booting this stick). Also all traditional sockets (PS2, serial, VGA) that we still need are available.
Our backup Vantage Pro Plus station was down since a couple of days, so I started with this to install the first T5730. This has been done yesterday: if you click on the "live-data" page to Backup -current what you see will come from the thin client (the humidity and solar irradiation sensors are dead; no repair planned in the near future).
The next replacement will be the main computer communicating with the datalogger. I ran 3 weeks of tests at home, and everything seems pretty reliable...
09 Apr 16 Cheers to Raoul who has managed to bring the system back to correct working... I do not yet have the details...

13:59 UTC: Raoul says the culprit was a flaky LAN patch cable... How bad that such small things can create as much havoc! (well, an atomic bomb also needs just a small neutron source to go off, and NASA mistaking inch for meters....)

08 Apr 16 Malfunction: The transfers from the datalogger to the other computers and the upload to the webserver are down. I am unable to restart from abroad... Raoul Tholl will check the problem asap... please be patient!

17:19 UTC: Raoul tried his best to bring back to proper working the datalogger computer... but it remains unreachable over the Lan. Possibly the Lan interface is defunct.... Watch the NEWS tomorrow.

01 Apr 16 Uploaded March 2016 (mar16.dat) data into data archive.
24 Mar 16 The pump that went bad tripped a fuse, which powered down the main switch of the meteoLCD LAN. This stopped transmission to the webserver yesterday from 11:30 on. "Repaired" today at 08:00.
This did not hold; NOx sensor shutdown until next maintenance by T.S.E.
23 Mar 16 The big external pump of the AC31M instrument (NO/NO2) was blocked (probably since 19/03). The inbuilt fan is broken, and the bearings seem in bad shape. As a quick repair, I de-gripped with MOS2 oil and added an external fan for cooling. Sensor restarted at 10:06 UTC. The pump will be replaced at the next major TSE overhaul in April.

16:00: The pump seems to be definitively broken. The NOx sensor will be shut down for the next couple of weeks.

20 Mar 16 The COM port problem showed up again 3 days ago; I replaced the computer by a spare yesterday, but no luck: problem again this morning!. As the problem is cured usually by a cold reboot (actually a power cycle of the computer), I inserted today ( at 16:45 UTC) a software reboot at every hour xx:15. Not sure if this helps....
12 Mar 16 The intermittent COM port problem has shown up again two times, the last yesterday. I have restarted the system and  will try to find a solution tomorrow.
05 Mar 16 The computer making the half-hour transfers from the datalogger and forwarding  the data to the Linux machine which creates the plots and uploads everything to the webserver, stopped working properly the 4th March at 04:00 UTC. The problem was that the old Win2k did not release the COM port used to access the datalogger. This is a rare problem, that happened before, and I am not sure about the real cause. It could be a timing problem in the AUTOMATE script which commands all the automatic stuff. Just to be sure I doubled a wait pause in the script... To bring back the computer to correct action one must power it down and up, a simple restart does not help. This was done today at 07:17 UTC.
01 Mar 16 1. I have changed the wrong date of all relevant files (today.dat, 7days.dat, all.dat), reset the logger's date to 01 Mar after a complete download, emptied the logger memory and restarted the logger at 08:00 UTC. As promised. no data have been lost... and the live plots are ok.
2. February data (feb16.dat) uploaded into data archive
3. Updated CO2 section in the "Trends" chapter; there were a couple of glitches (error in the regression lines equations)
29 Feb 16 Our old but reliable datalogger can not handle leap years; so the ASCII file you might download show today the date of 01th March. This problem will be handled tomorrow with a full download of the logger and a reset of the date to the correct day. With a bit of care no data will be lost.
06 Feb 16 Ultrasonic anemometer seems dubious. Plots will use old analog anemometer readings starting today 19:00 UTC
01 Feb 16 Upload of January data (jan16.dat) into data archive.
24 Jan 16 1. Calibration and maintenance work on all sensors
2.The trend analysis for 2015 is done. Please look here!
22 Jan 16 The NO/NO2 sensor AC31M is back and was restarted today at 10:00. The repair by T.S.E. was difficult and needed quite a lot of skill, as it had to be done with spare parts and at low cost, because meteoLCD would have been incapable to pay the price of a "standard" repair procedure. The peak in CO2 levels is not natural, but caused by human respiration during the installation and launching of the repaired sensor!
14 Jan 16 I finished a comparison between our total ozone column measurements made with the Microtops instruments and the Brewer #178 measurements made in Uccle. See the graph here. Our measurements are surprisingly good compared to the very big and expensive Brewer. Many thanks to Dr. Hugo de Backer, the director of, for providing the missing Uccle data series.
07 Jan 16 I started the update of the trends section today. Even with computers, this is a major work, so the update is an action in progress and will take some time to finish.
05 Jan
Upload of 2015_only.xls file with all 2015 data and statistics. This file contains also daily and monthly summaries and some comparisons with FINDEL data. Also upload of meteolcdstats_1998_2015.xls
02 Jan 16 1. Upload of dec15.dat into data archive
2. Created file of all monthly data files and added to archive.
01 Jan 16 1. The usual dogs-dinner in the 7days plot due to Gnuplot problems with handling previous and new year (the logger files do not contain the year number). This will be self healing in the the next 7 days.
2. Added overplot of Uccle and Diekirch total ozone column measurements for 2015 (see Dobson page)


Back Home